# Faster-than-Nyquist Signaling for Next-Generation Wireless: Principles, Applications, and Challenges

### v1

@@ II From Nyquist Signaling to FTN Signaling
where $x_{n}$ is the $n$-th entry of the transmitted symbol vector $\bf x$ of length $N$, $h\left(t\right)$ is a continuous energy-normalized TD pulse carrying the transmitted symbols, and $\tau T$ is the symbol duration. Particularly, $\tau\in\left(0,1\right]$ in (1) is the “acceleration factor” that controls the time separation between adjacent transmitted symbols. When $\tau=1$, (1) describes Nyquist signaling, while $\tau<1$, (1) represents FTN signaling. Some plots of signal with and without FTN appear in Fig. 1.
@@
$\triangleq$ $\left\{\begin{array}[]{cc}\sum\limits_{n=-\infty}^{\infty}\left|H\left(\frac{\xi}{\tau T}+n\right)\right|^{2},&{\rm if}\;\xi\in\left[-\frac{1}{2},\frac{1}{2}\right],\\ \quad\quad\quad\quad\quad\quad\quad\quad\;\;\;0,&\quad\quad\quad\quad\;\ {\rm else},\end{array}\right$. (4)
@@ IV-B Time-Domain Equalization
For both reduce-search and reduce-size equalization, employing an optimized filter after matched filtering can significantly improve error performance. For reduced-search BCJR, minimum-phase ISI models are typically preferred. By concentrating energy toward the early part of the ISI response, the trellis state selection becomes more accurate. Minimum-phase models can be constructed by applying all-pass filters to the received signal. In [Prlja2012MBCJR], a super minimum-phase model was constructed for FTN via such filtering, allowing $M$-BCJR to operate with $2$ to $4$ times fewer trellis states compared to the unfiltered model. For reduced-size BCJR, typically channel shortening filters are applied, which are designed for concentrating ISI energy into a reduced number of taps. These filters aim for reducing the effective channel memory so that neglecting the residual ISI only has minimal impact on error performance. Since FTN signaling typically has long but weak ISI tails, channel shortening has demonstrated promising error performance.
@@ IV-C Frequency-Domain Equalization
FDE is particularly attractive for FTN detection due to its low computational complexity, hence it has been adopted in diverse FTN studies. However, it also has its limitations. Firstly, the use of the CP imposes a throughput erosion compared to TD equalization, which does not require CP. More critically, the performance of FDE degrades as $\tau$ decreases. According to [kim2016properties], the eigenvalues of the FTN channel matrix correspond to samples of the folded-spectrum, taken at intervals of $\frac{1}{N\tau T}$, where $N$ denotes the number of transmitted symbols. Consequently, when $\tau<\frac{1}{WT}$, the folded-spectrum contains spectral nulls. In such cases, the corresponding subchannels carry no reliable information, hence resulting in severe performance degradation. To address this issue, FD power allocation has been considered. An eigenvalue decomposition (EVD)-based precoding scheme was introduced in [Ishihara2022Eigen], where a power allocation matrix is applied across the frequency bins to address the aforementioned problem. Specifically, when this power allocation follows the water-filling principle, the scheme can theoretically achieve the unconstrained channel capacity [zhang2025capacity]. As discussed in Section III, this approach is functionally equivalent to transmitting Nyquist signals using a sinc shaping pulse with an increased bandwidth.

### v2

@@ II From Nyquist Signaling to FTN Signaling
where $x_{n}$ is the $n$-th entry of the transmitted symbol vector $\bf x$ of length NN, $h\left(t\right)$ is a continuous energy-normalized TD pulse carrying the transmitted symbols, and $\tau T$ is the symbol duration. Particularly, $\tau\in\left(0,1\right]$ in (1) is the “acceleration factor” that controls the time separation between adjacent transmitted symbols. When $\tau=1$, (1) describes Nyquist signaling, while $\tau<1$, (1) represents FTN signaling. Some plots of signal with and without FTN appear in Fig. 1.
@@
$\triangleq$ $\left\{\begin{array}[]{cc}\sum\limits_{n=-\infty}^{\infty}\left|H\left(\frac{\xi+n}{\tau T}\right)\right|^{2},&{\rm if}\;\xi\in\left[-\frac{1}{2},\frac{1}{2}\right],\\ \quad\quad\quad\quad\quad\quad\;\;\;\;\;0,&\quad\quad\quad\quad\;\ {\rm else},\end{array}\right$. (4)
@@ IV-B Time-Domain Equalization
For both reduce-search and reduce-size equalization, employing an optimized filter after matched filtering can significantly improve error performance. For reduced-search BCJR, minimum-phase ISI models are typically preferred. By concentrating energy toward the early part of the ISI response, the trellis state selection becomes more accurate. Minimum-phase models can be constructed by applying all-pass filters to the received signal. In [Prlja2012MBCJR], a super minimum-phase model was constructed for FTN via such filtering, allowing $M$-BCJR to operate with 22 to 44 times fewer trellis states compared to the unfiltered model. For reduced-size BCJR, typically channel shortening filters are applied, which are designed for concentrating ISI energy into a reduced number of taps. These filters aim for reducing the effective channel memory so that neglecting the residual ISI only has minimal impact on error performance. Since FTN signaling typically has long but weak ISI tails, channel shortening has demonstrated promising error performance.
@@ IV-C Frequency-Domain Equalization
FDE is particularly attractive for FTN detection due to its low computational complexity, hence it has been adopted in diverse FTN studies. However, it also has its limitations. Firstly, the use of the CP imposes a throughput erosion compared to TD equalization, which does not require CP. More critically, the performance of FDE degrades as $\tau$ decreases. According to [kim2016properties], the eigenvalues of the FTN channel matrix correspond to samples of the folded-spectrum, taken at intervals of $\frac{1}{N\tau T}$, where NN denotes the number of transmitted symbols. Consequently, when $\tau<\frac{1}{WT}$, the folded-spectrum contains spectral nulls. In such cases, the corresponding subchannels carry no reliable information, hence resulting in severe performance degradation. To address this issue, FD power allocation has been considered. An eigenvalue decomposition (EVD)-based precoding scheme was introduced in [Ishihara2022Eigen], where a power allocation matrix is applied across the frequency bins to address the aforementioned problem. Specifically, when this power allocation follows the water-filling principle, the scheme can theoretically achieve the unconstrained channel capacity [zhang2025capacity]. As discussed in Section III, this approach is functionally equivalent to transmitting Nyquist signals using a sinc shaping pulse with an increased bandwidth.